Find the center (h,k) and radius r of the circle with the given equation.
4x^2+4y^2-12x+16y-5=0
Can you please break it down in simple steps?

Answer by ewatrrr(24785)

Re TY
the 5/4 came in two steps... see below (Generally these types do not involve a fraction)
4x^2+4y^2-12x+16y-5=0
4x^2 - 12x +4y^2 + 16y = 5 ****added 5 to both sides of the EQ
x^2 - 3x + y^2 + 4y = 5/4 ****divided thru by 4
.......
x^2 - 3x + (3/2)^2 + y^2 + 4y + 4 = 5/4 + (3/2)^2 + 4
(x- 3/2)^2 + (y+2)^2 = 5/4 + 9/4 + 16/4
.........
(x- 3/2)^2 + (y+2)^2 = 30/4